View Map
Map here
Overlay transparency


    War Diary: 211 Field Park Company RE

    Month and year: December 1944

    During the December 1944 War Diary for 211 Field Park Company, the unit spent the first 3 weeks of the month carrying out its static role being chiefely engaged on routine works but in the last 10 days moving to 3 different locations.

    National Archive Reference: WO 171/1591

    The "References to Appendices" column for this document was empty so has been omitted.

    All diaries for this unit:

    Please use the contact us page to report any errors or inaccuracies.

    Viewing map locations

    Many of the location in this diary have been geocoded, these locations are highlighted clicking on them will display them on a map. Many location also have historical maps and aerial photographs of the area available.

    « Return to war diaries

    Page 1 of 3
    PlaceDateHourSummary of Events and Information
    HOLLAND STEIN. 6436531- 4 Unit static. All sections performing normal functions. Nothing exceptional to report.
    5Chief Engineer visited unit and inspected Workshops and Field Stores. He interviewed C.Q.M.S.
    Larkman and recommended his application for a Quartermaters commission.
    5 - 10Still routine works. Hutting work being carried out under supervision of S/Sgt. Beaver R.E. at
    M.R. ______ completed a.m. 8. Unit allotted one days use of range at MR. 499628 - Coy revision
    on firing of rifle and bren and practised in firing of live P.I.A.T. Weather continues fine but cold.
    R.A.F. RD Squad ceased to be u/comd and returned to parent unit - this action welcomed on account of
    lack of bomb disposal work and consequent inability to keep squad fully employed. Since unit left
    FRANCE there has been practically no bomb disposal work except minor instances during stay at NIJMEGEN.
    11Detachment of some dozen tradesmen left to fit up 30 Corps H.Q. in secret location. 30 Corps Tps.
    R.E. come under command 12 Corps. All Corps flashes removed from tunics and painted off vehicles.
    12Very little to report during whole of this period. Unit static and normal routine works carried out.
    191600Received warning order from C.R.E. 30 Corps Tps. Engrs. to move. Adv. party to be prepared to move
    w.e.f. 1200 hrs 20. Following transport allocated to effect speedy move:- 15 GT 3 tonners of
    288 Coy. RASC. 30 GT lorries of 67 Coy RASC. followed by 50 vehicles of 128 Br. Coy. RASC. The two
    sections of Pioneers at present with this H.Q. to move forward with first lift of Engineer Stores
    and to assist in off loading arriving vehicles at new location. Two further sections of Pioneers in
    present area are allocated daily for loading. Rear party of 1 Offrs. 15 O.Rs. to remain to supervise
    clearance of Engr Stores, effect handover of Sawmill and any other outstanding matters.
    20No instructions during whole day. Coy standing by to move.
    2300Ordered to move 0100 hrs 21.
    BELGIUMBOURG LEOPOLD,210100Moved to BOURG LEOPOLD. Arrived 0730 hrs. Corps Tps. revert to under command 30 Corps. Leagured
    night 21/22.
    220900O.C. and Lt. WHITE carried out accommodation recces in area between LOUVAIN AND AERSCHOT, Intention
    to move today if possible. O.C. returned 0730 hrs - Coy to move to new location at LOUVAIN on 23.
    LOUVAIN J.87259623 240730Coy moved. Arrived LOUVAIN 0945 hrs.
    Coy settled in and prepared billets for occupation. 400 tons of Engineer Stores held in Field Stores.
    Weather extremely cold.
    Page 2 of 3
    PlaceDateHourSummary of Events and Information
    25No works today. Xmas dinner attended by C.R.E. 30 Corps Tps. (Lt. Col. R.E. BLACK. D.S.O., R.E.).
    Weather again cold. Leave ballot held today for U.K. leave. 42 Vacancies to this unit
    for Jan.
    26Routine works resumed. Some enemy air activity during night 26/27. Frosty weather continued.
    27Nothing to report. Routine works continued.
    28Informed that H.Q. 30 Corps Tps Engrs. and Fd Coys moving to new location 29. Ordered to provide
    maintenance party to assist in move of H.Q. 30 Corps (Main). 128 Br. Coy Adv. H.Q., 23 Mech Eqpt Pl.
    and 4 Dog Pl. to come under command on move of H.Q. 30 Corps Troops Engrs.
    291000O.C. visited C.E. 30 Corps and informed Coy to prepare to move in next two or three days.
    1200Maintenance party for H.Q. 30 Corps (Main) moved.
    1400O.C: visited C.R.E. 30 Corps Tps. Engrs in new location at JENEFFE MR. 3232 15 kilometer W of LIEGE.
    Routine works. Weather still very cold and roads icy,
    1800O.C. returned. Accommodation recces to go out a.m. 30 to area approx 10 miles N of NAMUR. Move if
    recce successful to take place 31. Shortage of transport for ferrying Field Stores necessitates
    leaving rear party consisting of whole of Field Stores and 2 Section of Pioneers to continue loading
    of vehicles engaged in ferrying to new location.
    300900Lt. BERRIDGE and Interpreter proceeding on accommodation recces. Coy preparing to move.
    1700Lt. Berrid ge returned. Suitable location found at EGHEZEE .872596. Coy to move 1100 hrs 31.
    EGHEZEE311100Coy moved. Reached new location 1400 hours. Coy billeted mainly in private houses. Weather
    continued exceptionally cold but fine. Rest of day spent settling in. 30 basic loads of Field
    Stores being ferried down. Remaining 300 tons left at LOUVAIN with small rear party.
    SUMMARY. For the first 3 weeks of this month the unit carried out its static role being chiefely
    engaged on routine works. In last ten days however unit has been in three different locations.
    It, at first, appeard that Xmas arrangements would have to be cancelled, but the short breathing
    space at LOUVAIN enabled unit to spend Xmas day in traditional atmosphere and spirit. Dinner was
    attended by C.R.E. 30 Corps Tps. Engrs. (Lt. Col. R.E. BLACK D.S.O., R.E.) and immediately prior
    to the dinner the first for ballot for leave in the U.K. was held. This units allotment for January 1945
    is 42.
    Sections have been kept moderately busy. M,D.R.S. dealt with 21 detectors, 19 of which were repaired
    Page 3 of 3
    PlaceDateHourSummary of Events and Information
    SUMMARY (conta).
    and 2 B.L.R'd. In this campaign to date 542 detectors have been received for repair. Total
    repaired 505. B.L.R'd 37. Total number of units served 58.
    Field Stores have been moderately busy with nothing exceptional to report.
    Workshops Section and 12 E & M Pl. working as combined team carried out the following works:-
    Making and painting of more than 1900 signs. 400 temporary wooden crosses, building wooden body
    on trailer, painting, construction of numerous carpentry jobs. e.g. wooden crate, dough trough
    (for Fd Bakery),filing boxes and cabinets etc. In addition to general maintenance at H.Q. 30
    Corps following were constructed - 1 map table 26' X 13', 6 map tables of miscellaneous sizes.
    4 map boards. A Jeep was winterproofed, partitions and blackouts made and fitted etc. Metal
    workers were engaged in routine work e.g. M.T. repairs, plant maintenance, construction of brackets
    etc. with considerable welding and turning work done. In addition preparation and maintenance
    of unit billets on moving and during occupation fell to these sections.